The nucleus of a 125Xe atom (an isotope of the element xenon with mass 125 u) is 6.0 fm in diameter. It has 54 protons and charge q=+54e (1 fm = 1 femtometer = 1×10−15m.) Hint: Treat the spherical nucleus as a point charge. What is the electric force on a proton 8.0 fm from the surface of the nucleus? Express your answer in newtons. What is the proton's acceleration?
Answer #1

distance between cnetr to proton, d = r + h

= (d/2) + h

= 6/2 + 8

= 11 fm

= 1.1*10^-14 m

a) F = k*Q*q/d^2

= 9*10^9*54*1.6*10^-19*1.6*10^-19/(1.1*10^-14)^2

= 102.8 N

b) a = F/m

= 102.8/(1.67*10^-27)

= 6.16 *10^28 m/s^2